# Low-Rank Gradient Compression with Error Feedback for MIMO Wireless Federated Learning

### v1

@@ Abstract
This paper presents a novel approach to enhance the communication efficiency of federated learning (FL) in multiple input and multiple output (MIMO) wireless systems. The proposed method centers on a low-rank matrix factorization strategy for local gradient compression based on alternating least squares, along with over-the-air computation and error feedback. The proposed protocol, termed over-the-air low-rank compression (Ota-LC), is demonstrated to have lower computation cost and lower communication overhead as compared to existing benchmarks while guaranteeing the same inference performance. As an example, when targeting a test accuracy of $80\%$ on the Cifar-10 dataset, Ota-LC achieves a reduction in total communication costs of at least $30\%$ when contrasted with benchmark schemes, while also reducing the computational complexity order by a factor equal to the sum of the dimension of the gradients.
@@ I Introduction
Federated Learning (FL) is a distributed machine learning paradigm designed to collaboratively train a shared model across multiple devices without the need to access their raw data [1]. The deployment of FL at the network edge introduces significant communication overhead due to the need to exchange high-dimensional model updates. This communication challenge becomes even more pronounced in the context of multiple-antennas wireless networks operating over shared radio resources.
@@
With noiseless, interference-free, but constrained communication, gradient compression has emerged as an efficient approach for reducing communication overhead. Representative gradient compression methods fall into three main categories, namely quantization, sparsification, and low-rank approximation. Gradient quantization supports digital communication, whereby each entry of the gradient is represented by using a limited number of bits [2], or even one-bit [3]. Gradient sparsification decreases the number of non-zero entries in the communicated gradient, with methods including top-k [4], rand-k [5], and threshold-based selection [6]. Compressive sensing (CS) recovery algorithms can further increase the sparsing level, supporting both analog and digital transmission [7, 8, 9]. Low-rank approximation, the focus of this work, uses low-rank factors to approximate the high dimensional gradient. These factors can be derived by projecting the original gradient into low-dimensional subspaces shared by multiple devices, resulting in random linear coding (RLC) [10], or by employing power iteration to factorize the gradient matrix, as demonstrated by powerSGD [11].
@@
Introducing wireless channels in the implementation of FL presents both novel design challenges and opportunities, particularly in the exploration of over-the-air aggregation in FL updates and in the utilization of the degrees of freedom inherent in multiple input multiple output (MIMO) communications. For a simplified wireless communication model with additive white Gaussian noise channels, reference [12] applies powerSGD [11] for FL updates, integrating over-the-air computation to enhance communication efficiency. In the context of decentralized FL, the authors of [13] consider a fading channel model and leverage RLC as a gradient compression method. To further optimize communication efficiency, reference [7] applied sparsification with a CS algorithm, namely turbo-CS [14], for gradient compression. For a massive number of receiving antennas, the paper [8] shows that utilizing the best linear unbiased estimator for parallel decoding of each local gradient outperforms over-the-air computation.
This work considers a fading wireless FL setting, proposing a low-rank gradient factorization method that factorizes the local gradients by adopting a decentralized alternating least squares, together with error feedback. The algorithm, termed over-the-air low-rank compression (Ota-LC) is shown via experiments to outperform existing methods in terms of communication cost, while also reducing the computation cost.
@@ II-1 Federated Gradient Descent
$F({\mathbf{W}})=\frac{1}{D}\sum_{k=1}^{K}D_{k}F_{k}({\mathbf{W}})=\frac{1}{D}\sum_{k=1}^{K}\sum_{({\mathbf{u}},v)\in\mathcal{D}_{k}}f({\mathbf{W}};{\mathbf{u}},v),$ (2)
@@
where $\eta$ denotes the learning rate. The updated parameter tensor (6) is sent back to all devices on the downlink. As in much of previous works such as [8, 7, 13], we assume that downlink communication is ideal, which is practically well justified when the edge server communicates with a less stringent power constraint than the devices. The steps in (4)-(6) are repeated until a convergence condition is met.
@@ II-2 Communication Model
III Low-Rank Gradient Compression with Error Feedback
In this section, we present the proposed over-the-air low-rank compression (Ota-LC) method. Ota-LC reduces the communication overhead by imposing a low-rank parameterization to subtensors of tensor ${\mathbf{W}}$, which is known to serve as a useful inductive bias in many learning problems [11, 16]. We also exploit the temporal correlations exhibited by gradients across iterative steps [17], as well as analog communication [18] and error feedback, with the latter aiming at a reduction of the bias of the low-rank compressed gradient [19].
III-A Overview
@@
To reduce the communication overhead, Ota-LC applies a low-rank matrix factorization separately to subtensors of the overall gradient tensor ${\mathbf{G}}^{(t)}$. The approach applies to a wide range of neural network architectures, including multilayer perceptrons, convolutional networks, graph neural networks, and transformers. In fact, the high-dimension tensor of these models’ parameters can be partitioned into sub-tensors that can be reshaped into matrices suitable for low-rank factorization. For example, the gradient of a CNN takes the form of a 4-D tensor $\in\mathbb{R}^{O\times I\times H\times W}$, where $O$ represents the number of output channels, $I$ is the number of input channels, and $H\times W$ denotes the kernel size. For CNNs, one way to obtain sub-tensors is by partitioning along the dimension of the output channels [11].
Low-rank compression is carried out in a distributed way by means of one step of a Jacobi SCA iteration [20] for a factorization problem. This step leverages temporal correlation across iterations by adopting an initialization based on the last iterate, and it also incorporates error feedback.
@@
At each iteration, the devices transmit the updated local factorization matrices through the wireless MIMO channel using analog modulation. The server recovers the sum of the local factorizations from the received signal via over-the-air computation to obtain the low-rank factorizations of the global gradient. Ota-LC is summarized in Algorithm 1, and a detailed explanation is provided next.
@@
4:      Server broadcasts ${\mathbf{P}}^{(t-1)}$, ${\mathbf{Q}}^{(t-1)}$ and ${\mathbf{B}}_{k}^{(t)}$ to each device $k$
@@
7:          Obtain the global gradient $\widehat{{\mathbf{G}}}^{(t-1)}$ via ${\mathbf{P}}^{(t-1)}[{\mathbf{Q}}^{(t-1)}]^{\top}$
@@
9:          Obtain gradient ${\mathbf{G}}_{k}^{(t)}=\frac{D_{k}}{D}\nabla F_{k}({\mathbf{W}}^{(t)})$ using gradient descent (4)
@@
11:          Update the low-rank factors $\overline{{\mathbf{P}}}_{k}^{(t)}$ and $\overline{{\mathbf{Q}}}_{k}^{(t)}$ using $\widetilde{{\mathbf{G}}}_{k}^{(t)}$ in (14)
12:         Obtain the transmitted signal ${\mathbf{S}}_{k}^{(t)}$
@@
17:      Server updates the low-rank factors ${\mathbf{P}}^{(t)}$ and ${\mathbf{Q}}^{(t)}$ via (15)
@@
III-B Low-Rank Gradient Compression
We start by describing low-rank compression (LC) for matrix factorization in the ideal case of a centralized system. We write, with some abuse of notation, as ${\mathbf{G}}$ any one of the subtensors in which the gradient is partitioned. We reshape such subtensor ${\mathbf{G}}$ into an $m\times n$ matrix, also denoted as ${\mathbf{G}}\in\mathbb{R}^{m\times n}$.
The high-dimensional gradient matrix ${\mathbf{G}}$ is represented by using low-rank matrices ${\mathbf{P}}\in\mathbb{R}^{m\times r}$ and ${\mathbf{Q}}\in\mathbb{R}^{n\times r}$ with rank $r\ll\min(n,m)$, as ${\mathbf{G}}={\mathbf{P}}{\mathbf{Q}}^{\top}$. Finding the optimal matrix factorization is formulated as the minimization [21]
$\min_{{\mathbf{P}},{\mathbf{Q}}}\frac{1}{2}\left\|{\mathbf{P}}{\mathbf{Q}}^{\top}-{\mathbf{G}}\right\|_{F}^{2}+\lambda(\|{\mathbf{P}}\|_{F}^{2}+\|{\mathbf{Q}}\|_{F}^{2}),$ (8)
where $\lambda>0$ is a regularization parameter, which controls the norms of factors ${\mathbf{P}}$ and ${\mathbf{Q}}$, and $\|\cdot\|_{F}$ denotes the Frobenius norm of a matrix. The problem (8) is biconvex [22], and a centralized algorithmic solution is given by Jacobi SCA [20, Algorithm 1].
To elaborate, given initialization ${\mathbf{P}}^{[0]}$ and ${\mathbf{Q}}^{[0]}$, for each round $s=1,2,\ldots,S$, Jacobi SCA solves two surrogate convex problems
$\min_{{\mathbf{P}}}\frac{1}{2}\left\|{\mathbf{P}}[{\mathbf{Q}}^{[s-1]}]^{\top}-{\mathbf{G}}\right\|_{F}^{2}+\lambda\|{\mathbf{P}}\|_{F}^{2},$ (9)
$\min_{{\mathbf{Q}}}\frac{1}{2}\left\|{\mathbf{P}}^{[s-1]}{\mathbf{Q}}^{\top}-{\mathbf{G}}\right\|_{F}^{2}+\lambda\|{\mathbf{Q}}\|_{F}^{2}$ (10)
in parallel, where ${\mathbf{P}}^{[s]}$ and ${\mathbf{Q}}^{[s]}$ represent the current iterates. Closed-form optimal solution for problems (9) and (10) are
$\overline{{\mathbf{P}}}^{[s]}$ $={\mathbf{G}}{\mathbf{Q}}^{[s-1]}{\left[({{\mathbf{Q}}^{[s-1]}})^{\top}{\mathbf{Q}}^{[s-1]}+\lambda{\mathbf{I}}\right]}^{-1},$ (11)
$\overline{{\mathbf{Q}}}^{[s]}$ $={\mathbf{G}}^{\top}{\mathbf{P}}^{[s-1]}{\left[({{\mathbf{P}}^{[s-1]}})^{\top}{\mathbf{P}}^{[s-1]}+\lambda{\mathbf{I}}\right]}^{-1},$
where ${\mathbf{I}}$ is the identity matrix. Then, the updated factors are
${\mathbf{P}}^{[s]}$ $={\mathbf{P}}^{[s-1]}+\beta^{[s]}(\overline{{\mathbf{P}}}^{[s]}-{\mathbf{P}}^{[s-1]}),$ (12)
${\mathbf{Q}}^{[s]}$ $={\mathbf{Q}}^{[s-1]}+\beta^{[s]}(\overline{{\mathbf{Q}}}^{[s]}-{\mathbf{Q}}^{[s-1]}),$
where $\beta^{[s]}\in(0,1]$ is the step size of the update (12). At the last round $S$, we obtain an approximation of global gradient as ${\mathbf{G}}\approx{\mathbf{P}}^{[S]}[{\mathbf{Q}}^{[S]}]^{\top}$. As a remark, when $\beta^{[s]}=1$, the update (12) is equivalent to the classical Jacobi best-response scheme, and its Gauss-Seidel counterpart update corresponds to alternating least squares (ALS) [23].
III-C Distributed LC
Consider now a distributed setting. In this case, by (5), the gradient matrix ${\mathbf{G}}$ is obtained as the sum ${\mathbf{G}}=\sum_{k=1}^{K}{\mathbf{G}}_{k}$, where matrix ${\mathbf{G}}_{k}$ is available at the $k$-th device. Plugging ${\mathbf{G}}=\sum_{k=1}^{K}{\mathbf{G}}_{k}$ into (11), we can write the update (11) as the sums
$\overline{{\mathbf{P}}}^{[s]}=\sum\nolimits_{k=1}^{K}\overline{{\mathbf{P}}}_{k}^{[s]},\quad\overline{{\mathbf{Q}}}^{[s]}=\sum\nolimits_{k=1}^{K}\overline{{\mathbf{Q}}}_{k}^{[s]},$ (13)
where the pair $\overline{{\mathbf{P}}}_{k}^{[s]}\in\mathbb{R}^{m\times r}$ and $\overline{{\mathbf{Q}}}_{k}^{[s]}\in\mathbb{R}^{n\times r}$ is evaluated at device $k$ with ${\mathbf{G}}_{k}$ in lieu of ${\mathbf{G}}$. This approach, however, requires a large communication overhead, as low-rank factors $\overline{{\mathbf{P}}}_{k}^{[s]}$ and $\overline{{\mathbf{Q}}}_{k}^{[s]}$ must be exchanged across the rounds $s=1,2,\ldots,S$ of the LC update in (12).
III-D Ota-LC: Protocol
To address this issue, at each round $t$ of federated gradient descent, Ota-LC produces updated low-rank factors ${\mathbf{P}}^{(t)}$ and ${\mathbf{Q}}^{(t)}$ by using the previous iterates ${\mathbf{P}}^{(t-1)}$ and ${\mathbf{Q}}^{(t-1)}$ as the starting point, for a single step of update (12), i.e., for $S=1$. More precisely, we consider a parallel version of the calculation (11) given by
for each round $t$, yielding the updated low-rank factor ${\mathbf{P}}^{(t)}$ and ${\mathbf{Q}}^{(t)}$
${\mathbf{Q}}^{(t)}$ $={\mathbf{Q}}^{(t-1)}+\beta^{(t)}(\overline{{\mathbf{Q}}}^{(t)}-{\mathbf{Q}}^{(t-1)}).$
This approach is motivated by the observation that gradients typically exhibit strong temporal correlations across iterations [11, 17].
III-E Ota-LC: Signal Design
In Ota-LC, the aggregation (13) is implemented using over-the-air computation to enhance spectrum efficiency. To start, we stack the obtained $\overline{{\mathbf{P}}}_{k}^{(t)}$ and $\overline{{\mathbf{Q}}}_{k}^{(t)}$ in (14) as the matrix ${\mathbf{R}}_{k}^{(t)}=[(\overline{{\mathbf{P}}}_{k}^{(t)})^{\top},(\overline{{\mathbf{Q}}}_{k}^{(t)})^{\top}]^{\top}\in\mathbb{R}^{(m+n)\times r}$ to be further processed for transmission. At each device $k$, the matrix ${\mathbf{R}}_{k}^{(t)}$ is then arranged into a complex matrix ${\mathbf{C}}_{k}^{(t)}\in\mathbb{C}^{\frac{m+n}{2}\times r}$ as in, e.g., [7, 8].

### v2

# Efficient Wireless Federated Learning via Low-Rank Gradient Factorization

@@ Abstract
This paper presents a novel gradient compression method for federated learning (FL) in wireless systems. The proposed method centers on a low-rank matrix factorization strategy for local gradient compression that is based on one iteration of a distributed Jacobi successive convex approximation (SCA) at each FL round. The low-rank approximation obtained at one round is used as a “warm start” initialization for Jacobi SCA in the next FL round. A new protocol termed over-the-air low-rank compression (Ota-LC) incorporating this gradient compression method with over-the-air computation and error feedback is shown to have lower computation cost and lower communication overhead, while guaranteeing the same inference performance, as compared with existing benchmarks. As an example, when targeting a test accuracy of $70\%$ on the Cifar-10 dataset, Ota-LC reduces total communication costs by at least $33\%$ compared to benchmark schemes.
@@ I Introduction
Federated Learning (FL) is a distributed machine learning paradigm designed to collaboratively train a shared model across multiple devices without the need to access their raw data [1]. The deployment of FL at the network edge introduces significant communication overhead due to the need to exchange high-dimensional model updates. This communication challenge becomes even more pronounced in the context of wireless networks operating over shared radio resources.
@@
With noiseless, interference-free, but constrained communication, gradient compression has emerged as an efficient approach for reducing communication overhead. Representative gradient compression methods fall into three main categories, namely quantization, sparsification, and low-rank approximation. Gradient quantization supports digital communication, whereby each entry of the gradient is represented by using a limited number of bits [2], or even only one bit [3, 4], and it can utilize a Grassmann manifold codebook [5] or vector quantization [6]. Gradient sparsification decreases the number of non-zero entries in the communicated gradient, with methods including Top-$K$, Rand-$K$, and threshold-based selection [7, 8, 9]. Compressive sensing (CS) recovery algorithms can further increase the sparsing level, supporting both analog and digital transmission [10, 11, 12]. Low-rank approximation, the focus of this work, uses low-rank factors to approximate the high dimensional gradient. These factors can be derived by projecting the original gradient into low-dimensional subspaces shared by multiple devices, resulting in random linear coding (RLC) [13], or by employing power iteration to factorize the gradient matrix, as demonstrated by powerSGD [14].
@@
Introducing wireless channels in the implementation of FL presents both novel design challenges and opportunities, particularly in the exploration of over-the-air aggregation in FL updates and in the utilization of the degrees of freedom inherent in multiple input multiple output (MIMO) communications. For a simplified wireless communication model with additive white Gaussian noise channels, [15] applies powerSGD [14] for FL updates, integrating over-the-air computation to enhance communication efficiency. Considering a fading channel model, the authors of [16] leverage RLC as a gradient compression method in the context of decentralized FL. To further optimize communication efficiency, reference [10] applied sparsification with a CS algorithm, namely turbo-CS [17], for gradient compression. Despite successfully reducing communication overhead in wireless FL systems, these works either impose high computational complexity on local devices as RLC or CS algorithms involving high-dimensional matrix operations or have to compromise learning performance, thus motivating our work.
This work proposes a gradient compression method induced by solving a low-rank matrix factorization problem at each round of FL. The method centers on Jacobi successive convex approximation (SCA) [18]. In each FL round, it utilizes the approximated low-rank gradient factors from the previous round as the “warm start” initialization for the next round and reduces to a single Jacobi SCA iteration. Applying this method to fading MIMO channels together with over-the-air computation and error feedback, the algorithm, termed over-the-air low-rank compression (Ota-LC), is shown via experiments to outperform existing methods in terms of communication overhead, while also reducing the computational complexity.
@@ II-1 Federated Gradient Descent
$F({\mathbf{W}})=\frac{1}{D}\sum\nolimits_{k=1}^{K}D_{k}F_{k}({\mathbf{W}}),$ (2)
@@
where $\eta$ denotes the learning rate. The updated parameter tensor (6) is sent back to all devices on the downlink. As in much of previous works such as [11, 10, 16], we assume that downlink communication is ideal, which is practically well justified when the edge server communicates with a less stringent power constraint than the devices. The steps in (4)-(6) are repeated until a convergence condition is met, e.g., until the condition $\|{\bf G}^{(t)}\|\leq\epsilon$ is satisfied for some value of $\epsilon>0$ or until approaching the maximum number of communication rounds.
@@ II-2 Communication Model
III Low-Rank Gradient Factorization in Federated Learning
@@
In this section, we present the proposed over-the-air low-rank compression (Ota-LC) method. Ota-LC reduces the communication overhead by imposing a low-rank factorization to subtensors of tensor ${\mathbf{G}}^{(t)}$, which is known to serve as a useful inductive bias in many learning problems [14]. The approach applies to a wide range of neural network architectures, including multilayer perceptrons, convolutional networks, graph neural networks, and transformers. In fact, the high-dimension tensor of these model parameters can be partitioned into sub-tensors that can be reshaped into matrices suitable for low-rank factorization. For example, the gradient of a CNN takes the form of a 4-D tensor $\in\mathbb{R}^{O\times I\times H\times W}$, where $O$ represents the number of output channels, $I$ is the number of input channels, and $H\times W$ denotes the kernel size. For CNNs, one way to obtain sub-tensors is by partitioning along the dimension of the output channels [14].
We start by proposing the design of distributed gradient factorization, followed by the implementation in MIMO systems [19] with error feedback aiming at a reduction of the bias of the low-rank compressed gradient [4].
III-A Ota-LC: Distributed LC for Gradient Factorization
We first formulate the gradient compression as a low-rank matrix factorization problem in a centralized system and then derive its distributed counterpart for FL. With some abuse of notation, we write ${\mathbf{G}}^{(t)}$ for any one of the subtensors in which the gradient is partitioned at round $t$. We reshape such subtensor into an $m\times n$ matrix, also denoted as ${\mathbf{G}}^{(t)}\in\mathbb{R}^{m\times n}$. The high-dimensional global gradient matrix ${\mathbf{G}}^{(t)}$ is represented by using low-rank matrices ${\mathbf{P}}^{(t)}\in\mathbb{R}^{m\times r}$ and ${\mathbf{Q}}^{(t)}\in\mathbb{R}^{n\times r}$ with rank $r\ll\min(n,m)$, as ${\mathbf{G}}^{(t)}={\mathbf{P}}^{(t)}[{\mathbf{Q}}^{(t)}]^{\top}$. Finding the optimal matrix factorization is formulated as the minimization [20]
$\min_{{\mathbf{P}}^{(t)},{\mathbf{Q}}^{(t)}}\frac{1}{2}\left\|{\mathbf{P}}^{(t)}[{\mathbf{Q}}^{(t)}]^{\top}-{\mathbf{G}}^{(t)}\right\|_{F}^{2}+\lambda(\|{\mathbf{P}}^{(t)}\|_{F}^{2}+\|{\mathbf{Q}}^{(t)}\|_{F}^{2}),$ (8)
where $\lambda>0$ is a regularization parameter, which controls the norms of factors ${\mathbf{P}}^{(t)}$ and ${\mathbf{Q}}^{(t)}$, and $\|\cdot\|_{F}$ denotes the Frobenius norm of a matrix. The problem (8) is biconvex [20], and a centralized algorithmic solution is given by multi-iteration Jacobi SCA [18, Algorithm 1].
To elaborate, we introduce $s$ denoting Jacobi SCA iteration. Given initialization ${\mathbf{P}}^{(t,0)}$ and ${\mathbf{Q}}^{(t,0)}$, for each iteration $s=1,2,\ldots,S$, Jacobi SCA solves two surrogate convex problems
$\min_{{\mathbf{P}}}\frac{1}{2}\left\|{\mathbf{P}}[{\mathbf{Q}}^{(t,s-1)}]^{\top}-{\mathbf{G}}^{(t)}\right\|_{F}^{2}+\lambda\|{\mathbf{P}}\|_{F}^{2},$ (9)
and
$\min_{{\mathbf{Q}}}\frac{1}{2}\left\|{\mathbf{P}}^{(t,s-1)}{\mathbf{Q}}^{\top}-{\mathbf{G}}^{(t)}\right\|_{F}^{2}+\lambda\|{\mathbf{Q}}\|_{F}^{2}$ (10)
in parallel. The closed-form optimal solution for problems (9) and (10) are
$\overline{{\mathbf{P}}}^{(t,s)}$ $={\mathbf{G}}^{(t)}{\mathbf{Q}}^{(t,s-1)}{\left[({{\mathbf{Q}}^{(t,s-1)}})^{\top}{\mathbf{Q}}^{(t,s-1)}+\lambda{\mathbf{I}}\right]}^{-1},$ (11)
$\overline{{\mathbf{Q}}}^{(t,s)}$ $=[{\mathbf{G}}^{(t)}]^{\top}{\mathbf{P}}^{(t,s-1)}{\left[({{\mathbf{P}}^{(t,s-1)}})^{\top}{\mathbf{P}}^{(t,s-1)}+\lambda{\mathbf{I}}\right]}^{-1},$
for updating ${\mathbf{P}}^{(t,s)}$ and ${\mathbf{Q}}^{(t,s)}$ as
${\mathbf{P}}^{(t,s)}$ $={\mathbf{P}}^{(t,s-1)}+\beta^{(t,s)}(\overline{{\mathbf{P}}}^{(t,s)}-{\mathbf{P}}^{(t,s-1)}),$ (12)
${\mathbf{Q}}^{(t,s)}$ $={\mathbf{Q}}^{(t,s-1)}+\beta^{(t,s)}(\overline{{\mathbf{Q}}}^{(t,s)}-{\mathbf{Q}}^{(t,s-1)}),$
where ${\mathbf{I}}$ in (11) is the identity matrix, and $\beta^{(t,s)}\in(0,1]$ is the step size of the update (12). At the last round $S$, we obtain an approximation of global gradient as ${\mathbf{G}}^{(t)}\approx{\mathbf{P}}^{(t,S)}[{\mathbf{Q}}^{(t,S)}]^{\top}$. As a remark, when $\beta^{(t,s)}=1$, the update (12) is equivalent to the classical Jacobi best-response scheme, and its Gauss-Seidel counterpart update corresponds to alternating least squares (ALS)[20].
Consider now a federated learning setting. In this case, by (5), the gradient matrix ${\mathbf{G}}^{(t)}$ is obtained as the sum ${\mathbf{G}}^{(t)}=\sum_{k=1}^{K}{\mathbf{G}}_{k}^{(t)}$, where matrix ${\mathbf{G}}_{k}^{(t)}$ is only available at the $k$-th device. Plugging ${\mathbf{G}}^{(t)}=\sum_{k=1}^{K}{\mathbf{G}}_{k}^{(t)}$ into (11), we can write $\overline{{\mathbf{P}}}^{(t,s)}$ and $\overline{{\mathbf{Q}}}^{(t,s)}$ in the update (12) as the sums
$\overline{{\mathbf{P}}}^{(t,s)}=\sum\nolimits_{k=1}^{K}\overline{{\mathbf{P}}}_{k}^{(t,s)},\quad\overline{{\mathbf{Q}}}^{(t,s)}=\sum\nolimits_{k=1}^{K}\overline{{\mathbf{Q}}}_{k}^{(t,s)},$ (13)
where the pair $\overline{{\mathbf{P}}}_{k}^{(t,s)}\in\mathbb{R}^{m\times r}$ and $\overline{{\mathbf{Q}}}_{k}^{(t,s)}\in\mathbb{R}^{n\times r}$ is evaluated at device $k$ with ${\mathbf{G}}_{k}^{(t)}$ in lieu of ${\mathbf{G}}^{(t)}$ via (11).
This approach, however, requires a large communication and computation overhead, as low-rank factors ${\mathbf{P}}_{k}^{(t,s)}$ and ${{\mathbf{Q}}}_{k}^{(t,s)}$ must be updated and exchanged across the rounds $s=1,2,\ldots,S$ of the LC update in (12) for each FL round $t$. Motivated by the observation that gradients typically exhibit strong temporal correlations across iterations [21], we address this issue by producing updated low-rank factors ${\mathbf{P}}^{(t)}$ and ${\mathbf{Q}}^{(t)}$ in a single update step (12), i.e., $S=1$, and utilizing the approximation obtained at the last FL round ${\mathbf{P}}^{(t-1)}$ and ${\mathbf{Q}}^{(t-1)}$ as “warm start” of the initial reference point. More precisely, the distributed version of the calculation (11) is designed as
for each round $t$, yielding the updated ${\mathbf{P}}^{(t)}$ and ${\mathbf{Q}}^{(t)}$ as
${\mathbf{Q}}^{(t)}$ $={\mathbf{Q}}^{(t-1)}+\beta^{(t)}(\overline{{\mathbf{Q}}}^{(t)}-{\mathbf{Q}}^{(t-1)}),$
where $\overline{{\mathbf{P}}}^{(t)}=\sum_{k=1}^{K}\overline{{\mathbf{P}}}_{k}^{(t)}$ and $\overline{{\mathbf{Q}}}^{(t)}=\sum_{k=1}^{K}\overline{{\mathbf{Q}}}_{k}^{(t)}$ as in (13).
III-B Ota-LC: Implementation in MIMO Wireless System
@@
At each iteration, the devices transmit the updated local factorization matrices (14) through the wireless MIMO channel using analog modulation. The server recovers the sum of the local factorizations from the received signal via over-the-air computation to obtain the low-rank factorizations of the global gradient. Ota-LC is summarized in Algorithm 1, and a detailed implementation is provided next.
@@
4:      Server broadcasts the low-rank factors ${\mathbf{P}}^{(t-1)}$, ${\mathbf{Q}}^{(t-1)}$ and ${\mathbf{B}}_{k}^{(t)}$ to each device $k$
@@
7:          Obtain the global gradient $\widehat{{\mathbf{G}}}^{(t-1)}={\mathbf{P}}^{(t-1)}[{\mathbf{Q}}^{(t-1)}]^{\top}$
@@
9:          Obtain gradient ${\mathbf{G}}_{k}^{(t)}$ using gradient descent (4)
@@
11:          Update the factors $\overline{{\mathbf{P}}}_{k}^{(t)}$ and $\overline{{\mathbf{Q}}}_{k}^{(t)}$ in (14)
12:         Obtain the transmitted signal ${\mathbf{S}}_{k}^{(t)}$ via (19)
@@
17:      Server updates ${\mathbf{P}}^{(t)}$ and ${\mathbf{Q}}^{(t)}$ via (15)
@@
III-B1 Signal Processing
In Ota-LC, the aggregation (13) is implemented using over-the-air computation to enhance spectrum efficiency. To start, we stack the obtained $\overline{{\mathbf{P}}}_{k}^{(t)}$ and $\overline{{\mathbf{Q}}}_{k}^{(t)}$ in (14) as the matrix
${\mathbf{R}}_{k}^{(t)}=[(\overline{{\mathbf{P}}}_{k}^{(t)})^{\top},(\overline{{\mathbf{Q}}}_{k}^{(t)})^{\top}]^{\top}\in\mathbb{R}^{(m+n)\times r}.$ (16)
At each device $k$, the matrix ${\mathbf{R}}_{k}^{(t)}$ in (16) is then arranged into a complex matrix ${\mathbf{C}}_{k}^{(t)}\in\mathbb{C}^{\frac{m+n}{2}\times r}$ as
${\mathbf{C}}_{k}^{(t)}[i,:]={\mathbf{R}}_{k}^{(t)}[i,:]+j{\mathbf{R}}_{k}^{(t)}[\frac{m+n}{2}+i,:],i=1,\cdots,\frac{m+n}{2}.$ (17)